OpenStudy (anonymous):

Find all solutions in the interval [0,2pi] cos^2x+2cosx+1=0

OpenStudy (anonymous):

call y = cos x. Solve the quadratic equation: y^2 + 2y + 1 = 0. Double root: y = -1. Cos x = y = -1 --> x = Pi. Check: X= Pi; cos x = -1; f(x) = 1 - 2 + 1 = 0 correct.